OpenStudy (k8lyn911):

Differentiate the function. f(x) = ln(25 sin^2(x)) I got f'(x)= ((25/2)(x-sin(x)cos(x))/(25 sin^2 (x)) Is that right? If it is, how do I simplify it?

OpenStudy (solomonzelman):

no it is not

geerky42 (geerky42):

you got denominator correct, but not numerator you have \(25\sin^2x = 25(\sin x)^2\), you can use chain rule here

OpenStudy (solomonzelman):

I would re-write 25sin²x as (5sinx)² and sub u² instead of 5sin²

OpenStudy (k8lyn911):

So would it be \[\frac{ 2u }{ u^2 } = \frac{ 2 (5 \sin x) }{ 25 \sin^2 x } = \frac{ 10 \sin x }{ 25 \sin^2 x }\] ?

geerky42 (geerky42):

\(\Large (u^2)' = 2u\cdot u'\)

geerky42 (geerky42):

you have \(\ln (u^2) = \dfrac{(u^2)'}{u^2}=\dfrac{2u\cdot u'}{u^2}\)

geerky42 (geerky42):

given that \(u = 5\sin x\)

OpenStudy (k8lyn911):

Is this right? \[\ln(u^2) = \frac{ (u^2)' }{ u^2 } = \frac{ 2u \times u' }{ u^2 } = \frac{ 2(5 \sin x)(5 \cos x) }{ 25 \sin^2 x } = \frac{ 50 \sin x \cos x }{ 25 \sin^2 x }\]

zepdrix (zepdrix):

Ugh the u's make it so confusing in my opinion. But yes, very good :) Looks correct.

OpenStudy (aum):

\(\Large f(x) = \ln(25 \sin^2(x)) = \ln (5\sin(x))^2 = \\ \Large 2*\ln(5\sin(x)) = 2\ln(5) + 2\ln(\sin(x)) \) \(\Large f'(x) = 2\frac{1}{\sin(x)}*(\cos(x)) = 2\cot(x)\)

geerky42 (geerky42):

@aum

OpenStudy (aum):

lol

OpenStudy (k8lyn911):

Haha xD Thanks so much, you guys! :) There are so many notes in my binder now that it's hard to find what you're looking for.

OpenStudy (aum):

You are welcome.